OpenStudy (anonymous):

28a/4a^3 (its a fraction). Choices: a) 7a^2 ; b) 7/a^2 ; c) 1/7a^2

OpenStudy (anonymous):

which part don't u understand? let me help you..

OpenStudy (anonymous):

Pretty sure its 7a^2, just making sure, plus the other possible answers made me a bit confused :P

OpenStudy (anonymous):

mmm, no.. i think it's not the answer ;o

OpenStudy (anonymous):

Aghh, I'm confused now >.<

OpenStudy (anonymous):

in my opinion : 28a/4a^3 = 28/4 * a/a^3 = 7 * 1/(a^-1 * a^3) = 7 * 1/(a^(-1+3)) = 7 * 1/a^2 = 7/a^2 cmiiw :)

OpenStudy (anonymous):

Omg that made more sense than I thought :P Thankyou sooo muchhh!! Cmiiw? :P